### v1

@@ 4.2 Proof of Theorem 1
Strongly convex case. Recalling the Lyapunov function (8) with
(33) $c_{x}=\frac{16\gamma\beta L}{n\left(1-\bar{\rho}_{W}\right)},\quad c_{y}=\frac{256\gamma^{3}\beta^{3}L\bar{\rho}_{W}}{n\left(1-\bar{\rho}_{W}\right)^{3}}.$
Then, with the help of Lemma 2, 5 and 6 and noticing that
@@
(34) $\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]$ $\leqslant\left(1-\max\left\{\frac{\mu\beta\gamma}{2},\frac{1-\bar{\rho}_{W}}{8}\right\}\right)\mathbb{E}\left[V_{\beta k}\right]$
$\quad+e_{1}\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]+e_{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+e_{3}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]$
@@
(35) $\gamma\leqslant\min\left\{\frac{1}{4\sqrt{2}\beta L},\frac{1-\bar{\rho}_{W}}{18\beta L\sqrt{\bar{\rho}_{W}}},\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{40\beta L\bar{\rho}_{W}}\right\},$
we obtain the convergence result for the strongly convex case as shown in (10).
@@
(36) $\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ $\leqslant\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\frac{\gamma^{2}\beta\sigma^{2}}{n}$
@@
(37) $\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]\leqslant\frac{2\left(\mathbb{E}\left[V_{\beta k}\right]-\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]\right)}{\gamma\beta}+\frac{2\gamma\sigma^{2}}{n}+\frac{3328\gamma^{2}\beta^{2}L\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}.$
@@
(38) $\left(1-\frac{768\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}^{2}}{\left(1-\bar{\rho}_{W}\right)^{4}}\right)\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
$\leqslant\frac{2\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{32\gamma^{2}\beta^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)^{3}K}+\frac{16n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}}\sigma^{2}+\frac{192n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}$
@@
(39) $\gamma\leqslant\min\left\{\frac{1}{4\beta L},\frac{1-\bar{\rho}_{W}}{14\beta L\sqrt{\bar{\rho}_{W}}},\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{40\beta L\bar{\rho}_{W}}\right\},$
@@
(40) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$ $\leqslant\frac{C_{0}}{\left(1-\bar{\rho}_{W}\right)K}+\frac{416n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}$
@@
(41) $C_{0}:=4\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]+\frac{64\gamma^{2}\beta^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)^{2}}.$
Then, by incorporating (40) into (28) in Lemma 3 and setting the initial consensus error in (41) to zero, we complete the proof.
@@ 5.1 Synthetic Example
(42) $\underset{x\in\mathbb{R}^{p}}{\min}f\left(x\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{=:f_{i}}{\underbrace{\left(\mathbb{E}_{v_{i}}\left[\left(h_{i}^{T}x-\nu_{i}\right)^{2}+\frac{\mu}{2}\left\|x\right\|^{2}\right]\right)}}},$
@@
In Fig. 4, we present the communication and computation complexity of FlexGT to achieve an accuracy of $\epsilon=10^{-3}$ with $\alpha=1,2,\dots,8$ and $\beta=1,2,\dots,8$. The experiment was run 100 times for each parameter setting to obtain an average complexity. The results demonstrate that, for a fixed number of communication steps $\alpha$, increasing the number of computation steps per round $\beta$ reduces the communication complexity but increases the computation complexity (moving towards the upper left). However, when $\beta$ is fixed and $\alpha$ is increased, the computation complexity decreases initially, while the communication complexity first decreases and then increases. It is worth noting that the Pareto-optimal solutions (bold red lines) are achieved according to the value of $\alpha$ suggested by Corollary 2. For smaller $\mu$, more communication and computation are required, as shown in the plot on the right. These observations validate our theoretical findings and illustrate the trade-off between communication and computation in our distributed optimization algorithms.
@@
Figure 4: Communication and computation complexity of FlexGT algorithm to achieve $\epsilon=10^{-3}$ accuracy with $\alpha=1,2,\dots,8$ and $\beta=1,2,\dots,8$ on synthetic data. Each node is in an exponential graph of $n=20$ nodes with 5 neighbors.
@@ 5.2 MNIST Dataset
Specifically, we evaluate the trained model on the full test dataset and compare the testing accuracy of FlexGT and Acc-FlexGT against baseline methods, including DSGD, DSGT, K-GT, and DFL. As shown in Fig. 6, with $\alpha=\beta=3$, when the data heterogeneity is high (top plots), FlexGT and Acc-FlexGT outperform the baseline methods in both communication and computation efficiency, with Acc-FlexGT achieving further performance improvements. Notably, the computation complexity of K-GT is worse due to simply skipping communications. These results illustrate the effectiveness of the proposed algorithms. Additionally, when data heterogeneity is relatively low (bottom plots), the gap between DSGD-based and GT-based algorithms becomes smaller, highlighting the robustness of GT in handling data heterogeneity.
@@ 6 Conclusions
In this paper, we proposed FlexGT, a flexible snapshot GT method, along with its accelerated variant, Acc-FlexGT, for solving distributed stochastic optimization problems in non-i.i.d. scenarios. Leveraging a unified convergence analysis framework for strongly convex, convex, and nonconvex objective functions, we prove that FlexGT and Acc-FlexGT achieve linear or sublinear convergence rates with the best-known communication and computation complexity in most settings. In particular, Acc-FlexGT attains optimal iteration complexity in the nonconvex case (up to a logarithmic factor) and improves existing results in the strongly convex case. These findings highlight the inherent trade-off between communication and computation, with Acc-FlexGT achieving a Pareto-optimal solution when prior knowledge of the graph is available.
@@
Appendix A Supporting Lemmas
This section presents several supporting lemmas for the proof of Theorem 1.
Lemma 4 (Bounding client divergence).
Suppose Assumptions 24 hold. Let the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$. Then, we get that for all $k\geqslant 0$
@@ Appendix A Supporting Lemmas
(44) $\mathbb{E}\left[\left\|\mathbf{x}_{\beta\left(k+1\right)}-\mathbf{1}\bar{x}_{\beta k}\right\|^{2}\right]$ $\leqslant\mathbb{E}\left[\left\|\left(\bar{W}-\mathbf{J}\right)\tilde{\mathbf{x}}_{\beta k}-\gamma\bar{W}\sum_{j=0}^{\beta-1}{\mathbf{y}_{\beta k+j}}\right\|^{2}\right]$
@@
(45) $\mathbb{E}\left[\left\|\mathbf{y}_{\beta k+t}\right\|^{2}\right]$ $\leqslant 2\mathbb{E}\left[\left\|\mathbf{y}_{\beta k}\right\|^{2}\right]+4n\sigma^{2}$
@@
Lemma 5 (Consensus error).
@@
(46) $\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+\frac{4\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{8n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sigma^{2}.$
Proof.
@@
(47) $\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
@@
(48) $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k+t}\right\|^{2}\right]\leqslant 2\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+4n\sigma^{2},$
and letting the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$, we complete the proof.
Lemma 6 (Tracking error).
@@
(49) $\gamma\leqslant\min\left\{\frac{1}{4\beta L},\frac{1-\bar{\rho}_{W}}{14\beta L\sqrt{\bar{\rho}_{W}}}\right\}.$
Then, we have that for all $k\geqslant 0$,
(50) $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{3+\bar{\rho}_{W}}{4}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{18\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
@@
(51) $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+2n\bar{\rho}_{W}\sigma^{2}$
@@
where we have used Young’s inequality and $\mathbb{E}\left[\nabla G_{\beta k}\right]=\nabla F_{\beta k}$. Then, noticing that
@@
$=2\bar{\rho}_{W}\mathbb{E}\left[\left<\left(\bar{W}-\mathbf{J}\right)\nabla G_{\beta k},\nabla F_{\beta k}-\nabla G_{\beta k}\right>\right]\leqslant 2\bar{\rho}_{W}n\sigma^{2},$
@@
$2\bar{\rho}_{W}\mathbb{E}\left[\left<\nabla F_{\beta\left(k+1\right)},\nabla F_{\beta k}-\nabla G_{\beta k}\right>\right]\leqslant\bar{\rho}_{W}\mathbb{E}\left[\left\|\nabla F_{\beta\left(k+1\right)}-\nabla F_{\beta k}\right\|^{2}\right]+\bar{\rho}_{W}n\sigma^{2},$
we get
(52) $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{6\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
@@
From the above lemma, it can be found that there is an interdependence between the consensus error and the GT error. To break the coupling between these two errors, we need the following lemmas. For convenience, we set the initial consensus error to zero, i.e., $x_{i,0}=x_{0}\in\mathbb{R}^{p},i\in\left[n\right]$.
Lemma 7 (Accumulated consensus error).
Suppose Assumptions 24 hold. Let the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$. Then, for all $K\geqslant 1$,
(53) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
$\leqslant\frac{2\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{16n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}}\sigma^{2}+\frac{8\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}.$
Proof.

### v2

@@ 4.2 Proof of Theorem 1
Strongly convex case. Recall the Lyapunov function (8) with the coefficients set as follows:
$c_{x}=\frac{16\gamma\beta L}{n\left(1-\bar{\rho}_{W}\right)},\quad c_{y}=\frac{256\gamma^{3}\beta^{3}L\bar{\rho}_{W}}{n\left(1-\bar{\rho}_{W}\right)^{3}}.$ (37)
Then, by substituting the results obtained from Lemmas 25 and 6 into the Lyapunov function (8), and noticing that
@@
$\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]$ $\leqslant\left(1-\max\left\{\frac{\mu\beta\gamma}{2},\frac{1-\bar{\rho}_{W}}{8}\right\}\right)\mathbb{E}\left[V_{\beta k}\right]$ (38)
$\quad+e_{1}\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]$
$\quad+e_{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+e_{3}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]$
@@
$\gamma\leqslant\min\left\{\frac{1}{4\sqrt{2}\beta L},\frac{1-\bar{\rho}_{W}}{18\beta L\sqrt{\bar{\rho}_{W}}},\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{40\beta L\bar{\rho}_{W}}\right\},$ (39)
we obtain the convergence result for strongly convex case as shown in (10).
@@
$\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ (40)
$\leqslant\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\frac{\gamma^{2}\beta\sigma^{2}}{n}$
@@
$\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]$ $\leqslant\frac{2\left(\mathbb{E}\left[V_{\beta k}\right]-\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]\right)}{\gamma\beta}$ (41)
$\quad+\frac{2\gamma\sigma^{2}}{n}+\frac{3328\gamma^{2}\beta^{2}L\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}.$
@@
$\left(1-\frac{768\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}^{2}}{\left(1-\bar{\rho}_{W}\right)^{4}}\right)\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$ (42)
$\leqslant\frac{2\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{32\gamma^{2}\beta^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)^{3}K}$
$\quad+\frac{16n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}}\sigma^{2}+\frac{192n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}$
@@
$\gamma\leqslant\min\left\{\frac{1}{4\beta L},\frac{1-\bar{\rho}_{W}}{14\beta L\sqrt{\bar{\rho}_{W}}},\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{40\beta L\bar{\rho}_{W}}\right\},$ (43)
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$ $\leqslant\frac{C_{0}}{\left(1-\bar{\rho}_{W}\right)K}+\frac{416n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}$ (44)
@@
$C_{0}:=4\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]+\frac{64\gamma^{2}\beta^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)^{2}}.$ (45)
Then, by incorporating (44) into (32) and setting the initial consensus error in (45) to zero, i.e., $x_{i,0}=x_{j,0},\forall i,j\in[n]$, we complete the proof.
4.3 Proof of Corollary 1
Denote by $\hat{\gamma}$ the obtained upper bound of the stepsize in (9) and
the constant part of the centralized error and the topology effect part in (10)–(12), respectively, both of which are controlled by the stepsize $\gamma$. Here, $\upsilon_{0}$ denotes either $\mathbb{E}\left[V_{0}\right]$ or $\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]$, depending on the specific case under consideration. Then, for the strongly convex case, tune the stepsize between
$\left\{\frac{2\ln\left(\max\left\{2,\beta^{2}\mu^{2}\hat{K}/\varUpsilon_{1},\beta^{3}\mu^{3}\hat{K}^{2}/\varUpsilon_{2}\right\}\right)}{\beta\mu\hat{K}},\,\,\hat{\gamma}\right\}$
to ensure that the centralized error and topology effect part match the linear part in terms of the number of iterations $\hat{K}$. Specifically, if $\hat{\gamma}$ is smaller, we have
$\mathbb{E}\left[V_{\beta k}\right]$ $\leqslant\exp\left(-\frac{\mu\left(1-\bar{\rho}_{W}\right)^{2}}{L}K\right)\mathbb{E}\left[V_{0}\right]$ (46)
$\quad+\left(\frac{\varUpsilon_{1}}{\beta^{2}\mu^{2}\hat{K}}+\frac{\varUpsilon_{2}}{\beta^{3}\mu^{3}\hat{K}^{2}}\right)\mathbb{E}\left[V_{0}\right]$
Otherwise, noticing that
$\exp\left(-\min\left\{\frac{\mu\beta\gamma}{2},\frac{1-\bar{\rho}_{W}}{8}\right\}K\right)\mathbb{E}\left[V_{0}\right]$ (47)
$\leqslant\exp\left(-\ln\left(\max\left\{2,\frac{\beta^{2}\mu^{2}\hat{K}}{\varUpsilon_{1}},\frac{\beta^{3}\mu^{3}\hat{K}^{2}}{\varUpsilon_{2}}\right\}\right)\right)\mathbb{E}\left[V_{0}\right]$
$\leqslant\left(\frac{\varUpsilon_{1}}{\beta^{2}\mu^{2}\hat{K}}+\frac{\varUpsilon_{2}}{\beta^{3}\mu^{3}\hat{K}^{2}}\right)\mathbb{E}\left[V_{0}\right],$
we get the same result as (46). Letting the right-hand side of (46) further smaller than or equal to $\epsilon$, we obtain the complexity result in (13).
Similarly, for the convex and nonconvex cases, tuning the stepsize between
$\left\{\frac{1}{\sqrt{\varUpsilon_{1}\hat{K}}},\,\,\frac{1}{\sqrt[3]{\varUpsilon_{2}\hat{K}}},\,\,\hat{\gamma}\right\}$,
we obtain the complexity results in (14) and (15).
4.4 Proof of Corollary 2
The proof is similar to that of Corollary 1. By Lemma 1, we have $1-\bar{\rho}_{W}\geqslant 1/2$, and the topology effect part $\varUpsilon_{2}$ can be dominated when $\bar{\rho}_{W}$ is sufficiently small. Specifically, let $\hat{\gamma}$ denote the obtained upper bound of the stepsize in (9), and tune the stepsize $\gamma$ between
$\left\{\frac{2\ln\left(\max\left\{2,\beta^{2}\mu^{2}\hat{K}/\varUpsilon_{1}\right\}\right)}{\beta\mu\hat{K}},\,\,\hat{\gamma}\right\}$
for the strongly convex case; and
$\left\{\frac{1}{\sqrt{\varUpsilon_{1}\hat{K}}},\,\,\hat{\gamma}\right\}$
@@ 5.1 Synthetic Example
$\underset{x\in\mathbb{R}^{p}}{\min}f\left(x\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{=:f_{i}}{\underbrace{\left(\mathbb{E}_{v_{i}}\left[\left(h_{i}^{T}x-\nu_{i}\right)^{2}+\frac{\mu}{2}\left\|x\right\|^{2}\right]\right)}}},$ (48)
@@
In Fig. 4, we present the communication and computational complexity of FlexGT to achieve an accuracy of $\epsilon=10^{-3}$ with $\alpha=1,2,\dots,8$ and $\beta=1,2,\dots,8$. The experiment was run 100 times for each parameter setting to obtain an average complexity. The results demonstrate that, for a fixed number of communication steps $\alpha$, increasing the number of computation steps per round $\beta$ reduces the communication complexity but increases the computational complexity (moving towards the upper left). However, when $\beta$ is fixed and $\alpha$ is increased, the computational complexity decreases initially, while the communication complexity first decreases and then increases. It is worth noting that the Pareto-optimal solutions (bold red lines) are achieved according to the value of $\alpha$ suggested by Corollary 2. For smaller $\mu$, more communication and computation are required, as shown in the plot on the right. These observations validate our theoretical findings and illustrate the trade-off between communication and computation in distributed optimization algorithms.
@@
Figure 4: Communication and computational complexity of FlexGT algorithm to achieve $\epsilon=10^{-3}$ accuracy with $\alpha=1,2,\dots,8$ and $\beta=1,2,\dots,8$ on synthetic data. Each node is in an exponential graph of $n=20$ nodes with 5 neighbors.
@@ 5.2 MNIST Dataset
Specifically, we evaluate the trained model on the full test dataset and compare the testing accuracy of FlexGT and Acc-FlexGT against baseline methods, including DSGD, DSGT, K-GT, and DFL. As shown in Fig. 6, with $\alpha=\beta=3$, when the data heterogeneity is high (top plots), FlexGT and Acc-FlexGT outperform the baseline methods in both communication and computational efficiency, with Acc-FlexGT achieving further performance improvements. Notably, the computational complexity of K-GT is worse due to simply skipping communications. These results illustrate the effectiveness of the proposed algorithms. Additionally, when data heterogeneity is relatively low (bottom plots), the gap between DSGD-based and GT-based algorithms becomes smaller, highlighting the robustness of GT in handling data heterogeneity.
@@ 6 Conclusions
In this paper, we have proposed FlexGT, a flexible snapshot GT method, along with its accelerated variant, Acc-FlexGT, for solving distributed stochastic optimization problems in non-i.i.d. scenarios. Leveraging a unified convergence analysis framework for strongly convex, convex, and nonconvex objective functions, we showed that FlexGT and Acc-FlexGT achieve linear or sublinear convergence rates with the best-known communication and computational complexity in most settings. In particular, Acc-FlexGT attains optimal iteration complexity and optimal communication complexity in the nonconvex case (up to a logarithmic factor) and improves existing results in the strongly convex case. These findings highlight the inherent trade-off between communication and computation, with Acc-FlexGT achieving a Pareto-optimal solution when connectivity of the graph is available.
@@
This section presents several supporting lemmas for the proof of Theorem 1. We begin by bounding the client divergence of the decision parameter between two communications.
Lemma 4 (Client divergence)
Suppose Assumptions 24 hold. Let the stepsize satisfy $\gamma\leqslant\frac{1}{8\beta L}$. Then, we have that for all $k\geqslant 0$
$\mathbb{E}\left[\left\|\mathbf{x}_{\beta\left(k+1\right)}-\mathbf{1}\bar{x}_{\beta k}\right\|^{2}\right]$ (49)
@@
$\mathbb{E}\left[\left\|\mathbf{x}_{\beta\left(k+1\right)}-\mathbf{1}\bar{x}_{\beta k}\right\|^{2}\right]$ (50)
$\leqslant\mathbb{E}\left[\left\|\left(\bar{W}-\mathbf{J}\right)\tilde{\mathbf{x}}_{\beta k}-\gamma\bar{W}\sum_{j=0}^{\beta-1}{\mathbf{y}_{\beta k+j}}\right\|^{2}\right]$
@@
$\mathbb{E}\left[\left\|\mathbf{y}_{\beta k+t}\right\|^{2}\right]$ $\leqslant 2\mathbb{E}\left[\left\|\mathbf{y}_{\beta k}\right\|^{2}\right]+4n\sigma^{2}$ (51)
@@
Then, we bound the consensus error of the decision variables and prove that it contracts at each round, as stated in the following lemma.
Lemma 5 (Consensus error)
@@
$\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$ (52)
$\quad+\frac{4\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{8n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sigma^{2}.$
Proof .2.
@@
$\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$ (53)
@@
$\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k+t}\right\|^{2}\right]\leqslant 2\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+4n\sigma^{2},$ (54)
and letting the stepsize $\gamma\leqslant\frac{1}{4\beta L}$, we complete the proof.
Based on Lemma 4, we can further bound the gradient tracking error with a similar contraction property.
Lemma 6 (GT error)
@@
$\gamma\leqslant\min\left\{\frac{1}{4\beta L},\frac{1-\bar{\rho}_{W}}{14\beta L\sqrt{\bar{\rho}_{W}}}\right\}.$ (55)
Then, we have, for all $k\geqslant 0$,
$\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ (56)
$\leqslant\frac{3+\bar{\rho}_{W}}{4}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{18\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
@@
$\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ (57)
$\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+2n\bar{\rho}_{W}\sigma^{2}$
@@
where we have used Young’s inequality and $\mathbb{E}\left[G_{\beta k}\right]=\nabla F_{\beta k}$. Then, noticing that
@@
$=2\bar{\rho}_{W}\mathbb{E}\left[\left<\left(\bar{W}-\mathbf{J}\right)G_{\beta k},\nabla F_{\beta k}-G_{\beta k}\right>\right]\leqslant 2\bar{\rho}_{W}n\sigma^{2},$
@@
[…]
$\leqslant\bar{\rho}_{W}\mathbb{E}\left[\left\|\nabla F_{\beta\left(k+1\right)}-\nabla F_{\beta k}\right\|^{2}\right]+\bar{\rho}_{W}n\sigma^{2},$
we further obtain
$\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ (58)
$\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{6\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
@@
From Lemmas 5 and 6, it can be found that there is an interdependence between the consensus error and the GT error. To decouple these two errors, we introduce Lemmas 7 and 8, which bound the accumulated consensus and gradient tracking errors, respectively. For convenience, we initialize the consensus error to zero, i.e., $x_{i,0}=x_{0}\in\mathbb{R}^{p},i\in\left[n\right]$.
Lemma 7 (Accumulated consensus error)
Suppose Assumptions 24 hold. Let the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$. Then, for all $K\geqslant 1$, we have
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$ (59)
$\leqslant\frac{2\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{16n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}}\sigma^{2}$
$\quad+\frac{8\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}.$
Proof .4.